You buy a computer directly from the manufacturer for R7332,00 and agree to repay it in equal installment over 3 years at the end of each month, starting one month from now. The interest rate is 10.7% per year, compounded monthly. How much will you pay in total?


1
Expert's answer
2021-05-07T14:24:15-0400

"PV=733200," No. of months in 3 years "n =12\times 3=36 months"


Interest compounded monthly "i =\dfrac{10.7}{12}=0.891" % "=0.00891"


Using present value formula-


"PV=PMT(\dfrac{1-(1+i)^{-n}}{i})

\\[9pt]733200=PMT(\dfrac{1-(1+0.00891)^{-36}}{0.00891}\\[9pt]733200=PMT\times 30.7\\[9pt]PMT=\dfrac{733200}{30.7}=23882.73"


Total amount we have to pay "= 23882.73\times 36=R859778.50"